# An active learning method for solving competitive multi-agent decision-making and control problems

### v3

@@ VI-D Comparison results
We now compare the performance of Algorithm 1 against those offered by TensCalc [43], a centralized state-of-the-art method for GNE computation. We remark that, unlike our approach, TensCalc requires a fully centralized scenario where the agents’ cost functions $J_{i}$ and local constraint sets $\mathcal{X}_{i}$ have to be known by the central entity. We thus consider four examples provided with the toolbox associated to TensCalc [^1] Available at https://github.com/hespanha/tenscalc, namely a zero-sum game and three different min-max problems, illustrated and described in Fig. 13.
@@ VII Conclusion and outlook
The proposed learning-based approach paves the way to numerous extensions in several (possibly complementary) directions. Looking at the specific problem of learning a GNE in GNEPs, for instance, given the tight connection between an equilibrium solution and a minimizer of the so-called Nikaido-Isoda function [47], one may design a tailored procedure along the line of Algorithm 1 to locally approximate such a function. Finally, we remark that problem (7) is based completely on exploiting the current predictors $\hat{f}_{i}$. In alternative to the initial random exploration phase, one could also investigate possible solutions for better exploring the decision set $\Omega$, as done for example in reinforcement learning [48] and surrogate-based global optimization [49], and analyze its potential benefits.
@@ A-A Technical proofs
Proof of Lemma 3.3: Let $\bar{f}:\mathbb{R}^{n}\times\mathbb{R}^{p}\to\bar{\mathbb{R}}$ be defined by $\bar{f}(x,\theta)\coloneqq f(x,\theta)+\iota_{\mathcal{X}}(x)$. Since $\mathcal{X}$ is a convex set and $f$ is a convex function, then $\bar{f}$ is also a convex function for each given $\theta\in\mathbb{R}^{p}$. By applying [36, Th. 2.6], it follows that $\mathcal{M}(\theta)=\textrm{argmin}_{y\in\mathbb{R}^{n}}~{}\bar{f}(y,\theta)$ is a convex set for all $\theta\in\mathbb{R}^{p}$.
@@
To show continuity of $f^{\star}$ at a generic $\tilde{\theta}\in\mathbb{R}^{p}$, we need to prove that $\lim_{k\rightarrow\infty}f^{\star}(\theta^{k})=f^{\star}(\tilde{\theta}),$ for any sequence $\{\theta^{k}\}_{k\in\mathbb{N}}$ such that $\lim_{k\rightarrow\infty}\theta^{k}=\tilde{\theta}$. Consider such a generic sequence and let $x^{k}\in\mathcal{M}(\theta^{k})$ and $\tilde{x}\in\mathcal{M}(\tilde{\theta})$. In view of Assumption 3.2, $\theta\mapsto f(x^{k},\theta)$ is also convex and differentiable w.r.t. $\theta$ and $\tilde{x}\in\mathcal{M}(\tilde{\theta})$, i.e., $\tilde{x}$ is a global minimizer of $f(\cdot,\tilde{\theta})$, and therefore we have that
@@
$\begin{split}f^{\star}(\theta^{k})=f(x^{k},\theta^{k})&\geq f(x^{k},\tilde{\theta})+\frac{\partial f(x^{k},\theta)}{\partial\theta}\Big{\rvert}_{\theta=\tilde{\theta}}^{\top}(\theta^{k}-\tilde{\theta})\\ &\geq f(\tilde{x},\tilde{\theta})+\frac{\partial f(x^{k},\tilde{\theta})}{\partial\theta}\Big{\rvert}_{\theta=\tilde{\theta}}^{\top}(\theta^{k}-\tilde{\theta})\\ &\geq f(\tilde{x},\tilde{\theta})-D(\tilde{\theta})\|\theta^{k}-\tilde{\theta}\|,\end{split}$
@@
where the last inequality holds since the partial derivative $\partial f(x,\theta)/\partial\theta$ evaluated in $\tilde{\theta}$ is bounded w.r.t. $x$, with bound $D(\tilde{\theta})\coloneqq\textrm{sup}_{x\in\mathcal{X}}\left|\tfrac{\partial f(x,\theta)}{\partial\theta}\Big{\rvert}_{\theta=\tilde{\theta}}\right|\in\mathbb{R}$. Therefore, we obtain:
$f(\tilde{x},\tilde{\theta})-D(\tilde{\theta})\|\theta^{k}-\tilde{\theta}\|\leq f^{\star}(\theta^{k})=f(x^{k},\theta^{k})\leq f(\tilde{x},\theta^{k})$ (17)
@@
where the last inequality follows by the optimality of $x^{k}$. Since $\theta\mapsto f(x,\theta)$ is continuous and $\lim_{k\rightarrow\infty}\theta^{k}=\tilde{\theta}$, we have that $\lim_{k\rightarrow\infty}f(\tilde{x},\theta^{k})=f(\tilde{x},\tilde{\theta})$, and the proof concludes by applying the squeeze theorem to (17). $\blacksquare$
@@
Proof of Lemma 3.6: Lower semicontinuity of $\bar{f}(x,\theta)$ readily follows by the continuity of $f(x,\theta)$ at any given point $(\tilde{x},\tilde{\theta})\in\mathbb{R}^{n}\times\mathbb{R}^{p}$, as
$\lim_{(x,\theta)\rightarrow(\tilde{x},\tilde{\theta})}\inf\bar{f}(x,\theta)$ $=\lim_{\varepsilon\rightarrow 0}\left[\inf_{(y,\xi)\in\mathcal{B}_{\varepsilon}(\tilde{x},\tilde{\theta})}f(y,\xi)\right]$
@@
=$=\left\{\begin{array}[]{lll}f(\tilde{x},\tilde{\theta})&\mbox{if}&\tilde{x}\in\mathcal{X}\\ +\infty&\mbox{if}&\tilde{x}\not\in\mathcal{X}\end{array}\right.$
@@
which in turn implies that $\lim_{(x,\theta)\rightarrow(\tilde{x},\tilde{\theta})}\inf\bar{f}(x,\theta)\geq\bar{f}(\tilde{x},\tilde{\theta})$ for all $(\tilde{x},\tilde{\theta})\in\mathbb{R}^{n}\times\mathbb{R}^{p}$. To show level-boundedness, note that $\mathcal{X}$ is bounded and nonempty, and therefore, for all $\alpha\in\mathbb{R}$, $\left\{x\in\mathbb{R}^{n}\ \left|\ \bar{f}(x,\theta)\leq\alpha\right.\right\}\subseteq\mathcal{X}$ for all $\theta\in\mathbb{R}^{p}$. Thus, by setting $\mathcal{S}=\mathcal{X}$ and $g=\bar{f}$ in Definition 3.5, the claim follows. $\blacksquare$
@@
Proof of Lemma 3.7: By [36, Th. 2.6], each vector $x^{k}$ in (5) is uniquely defined since $\tfrac{1}{2}\|\cdot\|_{2}^{2}$ is strictly convex, proper because $\mathcal{X}$ is nonempty and therefore, by Lemma 3.3, $\mathcal{M}(\theta^{k})$ is a convex set. Moreover, since $x\mapsto f(x,\theta)$ is continuous for all $\theta\in\mathbb{R}^{p}$ and $\mathcal{X}$ is compact, by Lemma 3.6 the function $\bar{f}(x,\theta)$, defined by $\bar{f}(x,\theta)=f(x,\theta)+\iota_{\mathcal{X}}(x)$, is lower semicontinuous w.r.t. $x$ and level-bounded in $x$ locally uniformly w.r.t. $\theta$. Still in view of Lemma 3.3, we also have that $\lim_{k\rightarrow\infty}f^{\star}(\theta^{k})=f^{\star}(\tilde{\theta})$. Thus, from [36, Th. 1.17(b)] it follows that the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by (5) is bounded and all its cluster points lie in $\mathcal{M}(\tilde{\theta})$, which is however a singleton and contains $\tilde{x}$ only.
@@
To start we note that, by combining the definition of limit [50, Def. 5.4] and the properties of the loss function stated in Standing Assumption 4.1, along with those of the affine surrogates $\hat{f}_{i}$ described in §IV-D, $\lim_{k\to\infty}x_{i}^{k}=\bar{x}_{i}$ and $\lim_{k\to\infty}\hat{\boldsymbol{x}}^{k}_{-i}=\bar{\boldsymbol{x}}_{-i}$ so that $(\bar{x}_{i},\bar{\boldsymbol{x}}_{-i})\in\Omega$ directly imply that $\lim_{k\to\infty}\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))=v>0$, for any $\tilde{\theta}_{i}\notin\mathcal{A}_{i}$, while for all $\tilde{\theta}_{i}\in\mathcal{A}_{i}$ instead, $\lim_{k\to\infty}\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))=0$. Since we are assuming that $\lim_{k\to\infty}\theta_{i}^{k}=\tilde{\theta}_{i}$, by considering a generic $\theta_{i}^{k}$ obtained from (6) we know that for any $\tilde{\varepsilon}>0$ there exists some $\tilde{M}\coloneqq\tilde{M}(\tilde{\varepsilon})\geq 0$ so that $|\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\theta^{k}_{i}))-v|\leq\tilde{\varepsilon}$, i.e., $-\tilde{\varepsilon}+v\leq\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\theta^{k}_{i}))\leq\tilde{\varepsilon}+v$, for all $k\geq\tilde{M}$. On the other hand, we also know from $\lim_{k\to\infty}\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))=0$ that for any $\bar{\varepsilon}>0$ there exists some $\bar{M}\coloneqq\bar{M}(\bar{\varepsilon})\geq 0$ so that $\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))\leq\bar{\varepsilon}$ for all $k\geq\bar{M}$.
@@
Proof of Proposition 4.4: Let us prove that $f(\boldsymbol{x},\theta)=\sum_{i\in\mathcal{N}}\|x_{i}-(\nu^{k+1}_{i})^{\top}\boldsymbol{x}_{-i}-c^{k+1}_{i}\|^{2}_{2}$ satisfies Assumption 3.1 and 3.2. Since $\Omega$ is a nonempty polytope, it is convex and bounded. Moreover, function $f$ is quadratic and positive semidefinite, and therefore continuous and convex w.r.t. $\boldsymbol{x}$, for all $\theta\in\mathbb{R}^{p}$. In addition, $f(\boldsymbol{x},\theta)$ can be expressed also as:
$f(\boldsymbol{x},\theta)$ $=\|A(\boldsymbol{x})\theta-b(\boldsymbol{x})\|_{2}^{2}$
@@
$\begin{split}\underset{\boldsymbol{x}\in\Omega}{\textrm{sup}}~{}\left|\frac{\partial f(\boldsymbol{x},\theta)}{\partial\theta}\right|&=\underset{\boldsymbol{x}\in\Omega}{\textrm{max}}~{}\left|2A(\boldsymbol{x})^{\top}A(\boldsymbol{x})\theta+A(\boldsymbol{x})^{\top}b(\boldsymbol{x})\right|\\ &=\underset{i=1,\ldots,q}{\textrm{max}}~{}\pm\{2A(v_{i})^{\top}A(v_{i})\theta+A(v_{i})^{\top}b(v_{i})\}\end{split}$
@@ A-B Derivation of Kalman filter equations in (12)
The updates (12) are obtained by applying linear Kalman filtering to the following linear time-varying system (we omit the index $i$ for simplicity of notation):
$\begin{array}[]{rcl}\theta^{k+1}&=&\theta^{k}+\xi^{k}\\ x^{k}&=&(\phi^{k})^{\top}\theta^{k}+\zeta^{k},\end{array}$
where $\xi^{k}$, $\zeta^{k}$ are zero-mean white noise terms with covariance $\beta I$ and 1, respectively. The Kalman filter equations are:
@@
Since
$P^{k+\nicefrac{{1}}{{2}}}\phi^{k}=P^{k}\phi^{k}-P^{k}\phi^{k}(\phi^{k})^{\top}P^{k}\phi^{k}/(1+(\phi^{k})^{\top}P^{k}\phi^{k})$
$~{}~{}~{}~{}~{}=P^{k}\phi^{k}-P^{k}\phi^{k}(1+(\phi^{k})^{\top}P^{k}\phi^{k}-1)/(1+(\phi^{k})^{\top}P^{k}\phi^{k})$
$~{}~{}~{}~{}~{}=P^{k}\phi^{k}/(1+(\phi^{k})^{\top}P^{k}\phi^{k})=M^{k},$
we obtain $\theta^{k+1}=\theta^{k}+P^{k}\phi^{k}/(1+(\phi^{k})^{\top}P^{k}\phi^{k})(x^{k}-(\phi^{k})^{\top}\theta^{k})$.

### v4

@@ VI-D Comparison results
We now compare the performance of Algorithm 1 against those offered by TensCalc [45], a centralized state-of-the-art method for GNE computation. We remark that, unlike our approach, TensCalc requires a fully centralized scenario where the agents’ cost functions $J_{i}$ and local constraint sets $\mathcal{X}_{i}$ have to be known by the central entity. We thus consider four examples provided with the toolbox associated to TensCalc [^2] Available at https://github.com/hespanha/tenscalc., namely a zero-sum game and three different min-max problems, illustrated and described in Fig. 13.
@@ VII Conclusion and outlook
The proposed approach paves the way to numerous extensions. Looking at the specific problem of learning a GNE in GNEPs, for instance, given the tight connection between an equilibrium solution and a minimizer of the Nikaido-Isoda function [49], one may design a tailored procedure along the line of Algorithm 1 to locally approximate such a function. Finally, we remark that problem (7) is based completely on exploiting the current predictors $\hat{f}_{i}$. In alternative to the initial random exploration phase, one could also investigate possible solutions for better exploring the decision set $\Omega$, as done for example in reinforcement learning [50] and surrogate-based global optimization [51], and analyze its potential benefits.
@@ A-A Technical proofs
Proof of Lemma 3.3: Let $\bar{r}:\mathbb{R}^{n}\times\mathbb{R}^{p}\to\bar{\mathbb{R}}$ be defined by $\bar{r}(x,\theta)\coloneqq r(x,\theta)+\iota_{\mathcal{X}}(x)$. Since $\mathcal{X}$ is a convex set and $r$ is a convex function, then $\bar{r}$ is also a convex function for each given $\theta\in\mathbb{R}^{p}$. By applying [36, Th. 2.6], it follows that $\mathcal{M}(\theta)=\textrm{argmin}_{y\in\mathbb{R}^{n}}~{}\bar{r}(y,\theta)$ is a convex set for all $\theta\in\mathbb{R}^{p}$.
@@
To show continuity of $r^{\star}$ at a generic $\tilde{\theta}\in\mathbb{R}^{p}$, we need to prove that $\lim_{k\rightarrow\infty}r^{\star}(\theta^{k})=r^{\star}(\tilde{\theta}),$ for any sequence $\{\theta^{k}\}_{k\in\mathbb{N}}$ such that $\lim_{k\rightarrow\infty}\theta^{k}=\tilde{\theta}$. Consider such a generic sequence and let $x^{k}\in\mathcal{M}(\theta^{k})$ and $\tilde{x}\in\mathcal{M}(\tilde{\theta})$. In view of Assumption 3.2, $\theta\mapsto r(x^{k},\theta)$ is also convex and differentiable w.r.t. $\theta$ and $\tilde{x}\in\mathcal{M}(\tilde{\theta})$, i.e., $\tilde{x}$ is a global minimizer of $r(\cdot,\tilde{\theta})$, and therefore we have that
@@
$\begin{split}r^{\star}(\theta^{k})=r(x^{k},\theta^{k})&\geq r(x^{k},\tilde{\theta})+\frac{\partial r(x^{k},\theta)}{\partial\theta}\Big{\rvert}_{\theta=\tilde{\theta}}^{\top}(\theta^{k}-\tilde{\theta})\\ &\geq r(\tilde{x},\tilde{\theta})+\frac{\partial r(x^{k},\tilde{\theta})}{\partial\theta}\Big{\rvert}_{\theta=\tilde{\theta}}^{\top}(\theta^{k}-\tilde{\theta})\\ &\geq r(\tilde{x},\tilde{\theta})-D(\tilde{\theta})\|\theta^{k}-\tilde{\theta}\|,\end{split}$
@@
where the last inequality holds since the partial derivative $\partial r(x,\theta)/\partial\theta$ evaluated in $\tilde{\theta}$ is bounded w.r.t. $x$, with bound $D(\tilde{\theta})\coloneqq\textrm{sup}_{x\in\mathcal{X}}\left|\tfrac{\partial r(x,\theta)}{\partial\theta}\Big{\rvert}_{\theta=\tilde{\theta}}\right|\in\mathbb{R}$. Therefore, we obtain:
$r(\tilde{x},\tilde{\theta})-D(\tilde{\theta})\|\theta^{k}-\tilde{\theta}\|\leq r^{\star}(\theta^{k})=r(x^{k},\theta^{k})\leq r(\tilde{x},\theta^{k})$ (17)
@@
where the last inequality follows by the optimality of $x^{k}$. Since $\theta\mapsto r(x,\theta)$ is continuous and $\lim_{k\rightarrow\infty}\theta^{k}=\tilde{\theta}$, we have that $\lim_{k\rightarrow\infty}r(\tilde{x},\theta^{k})=r(\tilde{x},\tilde{\theta})$, and the proof concludes by applying the squeeze theorem to (17). $\blacksquare$
@@
Proof of Lemma 3.6: Lower semicontinuity of $\bar{r}(x,\theta)$ readily follows by the continuity of $r(x,\theta)$ at any given point $(\tilde{x},\tilde{\theta})\in\mathbb{R}^{n}\times\mathbb{R}^{p}$, as
$\lim_{(x,\theta)\rightarrow(\tilde{x},\tilde{\theta})}\inf~{}\bar{r}(x,\theta)$ $=\lim_{\varepsilon\rightarrow 0}\left[\inf_{(y,\xi)\in\mathcal{B}_{\varepsilon}(\tilde{x},\tilde{\theta})}r(y,\xi)\right]$
@@
=$=\left\{\begin{array}[]{lll}r(\tilde{x},\tilde{\theta})&\mbox{if}&\tilde{x}\in\mathcal{X}\\ +\infty&\mbox{if}&\tilde{x}\not\in\mathcal{X}\end{array}\right.$
@@
which in turn implies that $\lim_{(x,\theta)\rightarrow(\tilde{x},\tilde{\theta})}\inf~{}\bar{r}(x,\theta)\geq\bar{r}(\tilde{x},\tilde{\theta})$ for all $(\tilde{x},\tilde{\theta})\in\mathbb{R}^{n}\times\mathbb{R}^{p}$. To show level-boundedness, note that $\mathcal{X}$ is bounded and nonempty, and therefore, for all $\alpha\in\mathbb{R}$, $\left\{x\in\mathbb{R}^{n}\ \left|\ \bar{r}(x,\theta)\leq\alpha\right.\right\}\subseteq\mathcal{X}$ for all $\theta\in\mathbb{R}^{p}$. Thus, by setting $\mathcal{S}=\mathcal{X}$ and $g=\bar{r}$ in Definition 3.5, the claim follows. $\blacksquare$
@@
Proof of Lemma 3.7: By [36, Th. 2.6], each vector $x^{k}$ in (5) is uniquely defined since $\tfrac{1}{2}\|\cdot\|_{2}^{2}$ is strictly convex, proper because $\mathcal{X}$ is nonempty and therefore, by Lemma 3.3, $\mathcal{M}(\theta^{k})$ is a convex set. Moreover, since $x\mapsto r(x,\theta)$ is continuous for all $\theta\in\mathbb{R}^{p}$ and $\mathcal{X}$ is compact, by Lemma 3.6 the function $\bar{r}(x,\theta)$, defined by $\bar{r}(x,\theta)=r(x,\theta)+\iota_{\mathcal{X}}(x)$, is lower semicontinuous w.r.t. $x$ and level-bounded in $x$ locally uniformly w.r.t. $\theta$. Still in view of Lemma 3.3, we also have that $\lim_{k\rightarrow\infty}r^{\star}(\theta^{k})=r^{\star}(\tilde{\theta})$. Thus, from [36, Th. 1.17(b)] it follows that the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by (5) is bounded and all its cluster points lie in $\mathcal{M}(\tilde{\theta})$, which is however a singleton and contains $\tilde{x}$ only.
@@
To start we note that, by combining the definition of limit [52, Def. 5.4] and the properties of the loss function stated in Standing Assumption 4.1, along with those of the affine surrogates $\hat{f}_{i}$ described in §IV-D, $\lim_{k\to\infty}x_{i}^{k}=\bar{x}_{i}$ and $\lim_{k\to\infty}\hat{\boldsymbol{x}}^{k}_{-i}=\bar{\boldsymbol{x}}_{-i}$ so that $(\bar{x}_{i},\bar{\boldsymbol{x}}_{-i})\in\Omega$ directly imply that $\lim_{k\to\infty}\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))=v>0$, for any $\tilde{\theta}_{i}\notin\mathcal{A}_{i}$, while for all $\tilde{\theta}_{i}\in\mathcal{A}_{i}$ instead, $\lim_{k\to\infty}\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))=0$. Since we are assuming that $\lim_{k\to\infty}\theta_{i}^{k}=\tilde{\theta}_{i}$, by considering a generic $\theta_{i}^{k}$ obtained from (6) we know that for any $\tilde{\varepsilon}>0$ there exists some $\tilde{M}\coloneqq\tilde{M}(\tilde{\varepsilon})\geq 0$ so that $|\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\theta^{k}_{i}))-v|\leq\tilde{\varepsilon}$, i.e., $-\tilde{\varepsilon}+v\leq\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\theta^{k}_{i}))\leq\tilde{\varepsilon}+v$, for all $k\geq\tilde{M}$. On the other hand, we also know from $\lim_{k\to\infty}\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))=0$ that for any $\bar{\varepsilon}>0$ there exists some $\bar{M}\coloneqq\bar{M}(\bar{\varepsilon})\geq 0$ so that $\ell_{i}(x_{i}^{k},\hat{f}_{i}(\hat{\boldsymbol{x}}^{k}_{-i},\tilde{\theta}_{i}))\leq\bar{\varepsilon}$ for all $k\geq\bar{M}$.
@@
Proof of Proposition 4.4: Let us prove that $r(\boldsymbol{x},\theta)=\sum_{i\in\mathcal{N}}\|x_{i}-(\nu^{k+1}_{i})^{\top}\boldsymbol{x}_{-i}-c^{k+1}_{i}\|^{2}_{2}$ satisfies Assumption 3.1 and 3.2. Since $\Omega$ is a nonempty polytope, it is convex and bounded. Moreover, function $r$ is quadratic and positive semidefinite, and therefore continuous and convex w.r.t. $\boldsymbol{x}$, for all $\theta\in\mathbb{R}^{p}$. In addition, $r(\boldsymbol{x},\theta)$ can be expressed also as:
$r(\boldsymbol{x},\theta)$ $=\|A(\boldsymbol{x})\theta-b(\boldsymbol{x})\|_{2}^{2}$
@@
$\begin{split}\underset{\boldsymbol{x}\in\Omega}{\textrm{sup}}~{}\left|\frac{\partial r(\boldsymbol{x},\theta)}{\partial\theta}\right|&=\underset{\boldsymbol{x}\in\Omega}{\textrm{max}}~{}\left|2A(\boldsymbol{x})^{\top}A(\boldsymbol{x})\theta+A(\boldsymbol{x})^{\top}b(\boldsymbol{x})\right|\\ &=\underset{i=1,\ldots,q}{\textrm{max}}~{}\pm\{2A(v_{i})^{\top}A(v_{i})\theta+A(v_{i})^{\top}b(v_{i})\}\end{split}$
@@ A-B Derivation of Kalman filter equations in (12)
The updates (12) are obtained by applying linear Kalman filtering to the linear time-varying system (we omit the index $i$ for simplicity of notation): $\theta^{k+1}=\theta^{k}+\xi^{k}$, $x^{k}=(\phi^{k})^{\top}\theta^{k}+\zeta^{k}$, where $\xi^{k}$, $\zeta^{k}$ are zero-mean white noise terms with covariance $\beta I$ and 1, respectively. The Kalman filter equations are:
@@
Since $P^{k+\nicefrac{{1}}{{2}}}\phi^{k}=P^{k}\phi^{k}-P^{k}\phi^{k}(\phi^{k})^{\top}P^{k}\phi^{k}/(1+(\phi^{k})^{\top}P^{k}\phi^{k})=P^{k}\phi^{k}/(1+(\phi^{k})^{\top}P^{k}\phi^{k})=M^{k}$, we obtain $\theta^{k+1}=\theta^{k}+P^{k}\phi^{k}/(1+(\phi^{k})^{\top}P^{k}\phi^{k})(x^{k}-(\phi^{k})^{\top}\theta^{k})$.